OpenStudy (anonymous):

Pleaseeeee help How many liters of 75% solution should be mixed with a 55% solution to get 70 liters of 63% solution? How many liters of the 55% and 75% solutions should be used?

OpenStudy (anonymous):

first you need 63% of 70 which is \(.63\times 70=44.1\)

OpenStudy (anonymous):

then put \(x\) as the amount of 75% solution, making \(70-x\) the amount of 55% solution and solve \[.75x+.55(70-x)=44.1\] for \(x\)

OpenStudy (anonymous):

probably easier to get rid of the decimals and solve \[75x+55(70-x)=4410\]